what should be tension applied to a wire of length 1m and mass 10gm, if it has to vibrate with fundamental frequency of 50 Hz

Grade:12

2 Answers

Phanindra
6 years ago
n=(1/2l)ГT/Гμ
Here μ=mass(kg/length in meters i.e Linear mass density)
T = tension in the string
n= frequency and l = length of the string
Subtitute the values to get the value
T = 4(l×l)(n×n)×μ....
 
Arun
6 years ago
Dear student
mass m = 10 gm = 0.01
n = (1/2L) \sqrt(T/m)
Now
50 = 1/2 \sqrt(T/0.01)
100² = T/0.01
10000*0.01 = T
T = 100 N
Regards
Arun (askIITians forum expert)